# Does Refusal Training in LLMs Generalize to the Past Tense?

### v3

@@ 2 Methodology
Target LLMs. We evaluate eight target LLMs: Llama-3 8B (AI@Meta, 2024), Claude-3.5 Sonnet (Anthropic, 2024), GPT-3.5 Turbo (OpenAI, 2023), Gemma-2 9B (DeepMind, 2024), Phi-3-Mini (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024c), GPT-4o (OpenAI, 2024d), and R2D2 (Mazeika et al., 2024). Most of these models use supervised fine-tuning and reinforcement learning from human feedback for refusal training. In addition, R2D2 uses adversarial training against the GCG attack on top of SFT and DPO used to fine-tune the Zephyr model (Tunstall et al., 2023). For Llama-3 8B, we use the refusal-enhancing prompt introduced in Llama-2 (Touvron et al., 2023), while for the rest of the LLMs, we use their default system prompts. We list all system prompts in Appendix A.
@@ 3 Systematic Evaluation of the Past Tense Attack
4 Does Fine-tuning on the Past Tense Examples Help?
@@
Results. We show systematic results in Table 4, which suggest that it is straightforward to reduce the attack success rate to 0%. The overrefusal rate predictably increases with a higher proportion of refusal data in the fine-tuning mix. To provide some point of reference, the overrefusal rate of Llama-3 8B is 19%, while the ASR is 27% according to the GPT-4 judge. Thus, FT 2%/98% with 6% overrefusal rate and 24% ASR improves the Pareto frontier between correct and wrong refusals. We note that with more data, this trade-off could likely be improved further. Overall, defending against past-tense reformulations is feasible if one directly adds the corresponding data during fine-tuning, although wrong refusals must be carefully controlled.
@@
5 Are Reasoning Models More Robust?
Very recently, LLMs with advanced chain-of-thought reasoning capabilities were introduced, such as the o1-mini and o1-preview models (OpenAI, 2024b). Before outputting any answer, these models produce a long reasoning chain about how to fulfill a user request. Relevant to our work, the reasoning chains are also trained on their usage policies to teach human values and improve safety. They claim that such chain-of-thought reasoning achieves substantial improvements on their hardest jailbreak evaluations.
@@ 5 Are Reasoning Models More Robust?
Discussion. We notice the following pattern in successful jailbreaks. First, the summary of the internal reasoning acknowledges that a past tense request might be harmful. However, since it is a historical reference, the model often goes on to generating an answer, and sometimes reveals more detailed information that it should. It is also interesting to note that chain-of-thought based answers work like an implicit output filter due to a natural breakpoint between a chain of thought reasoning and actual answer. The presence of this breakpoint is an interesting departure from end-to-end autoregressive models, such as those presented in Section 3. Overall, reasoning models provide an interesting approach for improving safety that—in case of the o1 models—does not lead to significantly reduced refusal rate, but rather to less informative answers. We believe that the usefulness of the generated answers should be evaluated in a more nuanced way than is typically done in the current literature.
@@ 6 Related Work
Generalization of LLM alignment. After pretraining, LLMs are typically aligned to human preferences using techniques like SFT (Chung et al., 2022), RLHF (Ouyang et al., 2022), or DPO (Rafailov et al., 2023). One of the objectives of the alignment process is to make LLMs produce refusals on harmful queries, which involves adding refusal examples to the fine-tuning data. Since it is impossible to add all possible reformulations of harmful requests in the fine-tuning set, LLMs alignment crucially relies on the ability to generalize from a few examples per harmful behavior. Empirical studies support this capability: Dang et al. (2024) observe that RLHF generalizes from English to other languages, and Li et al. (2024b) make the same claim specifically for refusal training. This observation is consistent with Wendler et al. (2024) who argue that LLMs pretrained primarily on English data tend to internally map other languages to English. Therefore, fine-tuning on English data can suffice since the internal representations largely coincide with the representations of other languages. However, this capacity is in stark contrast to past-tense reformulations, which, as we show, represent a blind spot. We hypothesize that the underlying reason is that the internal representations for the past and present tenses are distinct. Thus, one has to include reformulations in both tenses to achieve more robust refusals. On a related note, Berglund et al. (2024) discuss the reversal curse phenomenon, i.e., training on ”A is B” does not make the model automatically generalize to ”B is A”. This failure again illustrates that generalization mechanisms taken for granted in humans do not necessarily occur in LLMs.
@@ 7 Discussion
We believe the main reason for this generalization gap is that past tense examples are out-of-distribution compared to the refusal examples used for fine-tuning, and current alignment techniques do not automatically generalize to them. Indeed, as we have shown in Section 4, correctly refusing on past tense examples is feasible via direct fine-tuning, and some models—like Llama-3 with the refusal-enhancing system prompt—are already relatively robust. Moreover, there are also other possible solutions that do not rely on SFT or RLHF, such as output-based detectors (Inan et al., 2023) and representation-based methods, including harmfulness probing and representation rerouting (Zou et al., 2024). These approaches can reject harmful outputs, which seems to be an easier task compared to patching all possible inputs that can lead to harmful generations.
@@ Acknowledgements
We thank Francesco Croce for many insightful discussions and feedback on a draft of this paper. We are grateful to OpenAI for providing us API credits as part of the OpenAI Researcher Access Program. We thank Sudarsh Kunnavakkam for providing the API key to run the experiments on Claude-3.5 Sonnet. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship. This work was supported by the Swiss National Science Foundation (grant number 212111).
@@ Appendix A Additional Details
Table 9: The system prompt for R2D2 from HarmBench (Mazeika et al., 2024).
@@
Table 11: The system prompt for GPT-4 as a semantic judge introduced in Chao et al. (2023).
@@
Table 12: The prompt for Llama-3-70B as a jailbreak judge from JailbreakBench (Chao et al., 2024).
@@
Table 13: The prompt for Llama-3 8B as a refusal judge from JailbreakBench (Chao et al., 2024).

### v4

@@ 2 Methodology
Target LLMs. We evaluate eight target LLMs: Llama-3 8B Instruct (AI@Meta, 2024), Claude-3.5 Sonnet (Anthropic, 2024), GPT-3.5 Turbo (OpenAI, 2023), Gemma-2 9B Instruct (DeepMind, 2024), Phi-3-Mini Instruct (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024c), GPT-4o (OpenAI, 2024d), and R2D2 (Mazeika et al., 2024). Most of these models use supervised fine-tuning and reinforcement learning from human feedback for refusal training. In addition, R2D2 uses adversarial training against the GCG attack on top of SFT and DPO used to fine-tune the Zephyr model (Tunstall et al., 2023). For Llama-3 8B, we use the refusal-enhancing prompt introduced in Llama-2 (Touvron et al., 2023), while for the rest of the LLMs, we use their default system prompts. We list all system prompts in Appendix A.
@@ 3 Systematic Evaluation of the Past Tense Attack
4 Does Fine-tuning on Past Tense Examples Help?
In this section, we explore whether direct fine-tuning on harmful past tense requests improves an LLM’s robustness to such requests.
@@ 4 Does Fine-tuning on Past Tense Examples Help?
Results. We show systematic results in Table 4, which suggest that it is straightforward to reduce the attack success rate to 0%. The overrefusal rate predictably increases with a higher proportion of refusal data in the fine-tuning mix. To provide some point of reference, the overrefusal rate of Llama-3 8B is 19%, while the ASR is 27% according to the GPT-4 judge. Thus, FT 2%/98% with 6% overrefusal rate and 24% ASR improves the Pareto frontier between correct and wrong refusals. We note that with more data, this trade-off could likely be improved further. Overall, defending specifically against past-tense reformulations is feasible if one directly adds the corresponding data during fine-tuning, although wrong refusals must be carefully controlled.
@@
5 Are Reasoning Models More Robust?
Very recently, LLMs with advanced chain-of-thought reasoning capabilities (Wei et al., 2022) were introduced, such as the o1-mini and o1-preview models (OpenAI, 2024b). Before outputting any answer, these models produce a long reasoning chain about how to fulfill a user request. Relevant to our work, the reasoning chains are also trained on their usage policies to teach human values and improve safety. They claim that such chain-of-thought reasoning achieves substantial improvements on their hardest jailbreak evaluations.
@@ 5 Are Reasoning Models More Robust?
Discussion. We notice the following pattern in successful jailbreaks. First, the summary of the internal reasoning acknowledges that a past tense request might be harmful. However, since it is a historical reference, the model often goes on to generating an answer, and sometimes reveals more detailed information that it should. It is also interesting to note that chain-of-thought based answers work like an implicit output filter due to a natural breakpoint between a chain-of-thought reasoning and actual answer. The presence of this breakpoint is an interesting departure from end-to-end autoregressive models, such as those presented in Section 3. Overall, reasoning models provide an interesting approach for improving safety that—in case of the o1 models—does not lead to significantly reduced refusal rate, but rather to less informative answers. We believe that the usefulness of the generated answers should be evaluated in a more nuanced way than is typically done in the current literature.
@@ 6 Related Work
Generalization of LLM alignment. After pretraining, LLMs are typically aligned to human preferences using techniques like SFT (Chung et al., 2022), RLHF (Ouyang et al., 2022), or DPO (Rafailov et al., 2023). One of the objectives of the alignment process is to make LLMs produce refusals on harmful queries, which involves adding refusal examples to the fine-tuning data. Since it is impossible to add all possible reformulations of harmful requests in the fine-tuning set, LLMs alignment crucially relies on the ability to generalize from a few examples per harmful behavior. Empirical studies support this capability: Dang et al. (2024) observe that RLHF generalizes from English to other languages, and Li et al. (2024b) make the same claim specifically for refusal training. This observation is consistent with Wendler et al. (2024) who argue that LLMs pretrained primarily on English data tend to internally map other languages to English. Therefore, fine-tuning on English data can suffice since the internal representations largely coincide with the representations of other languages. However, this capacity is in stark contrast to past-tense reformulations, which, as we show, represent a blind spot. We hypothesize that the underlying reason is that the internal representations for the past and present tenses are distinct. Thus, as illustrated in Section 4, one has to include reformulations in both tenses to the training set to achieve more robust refusals. On a related note, Berglund et al. (2024) discuss the reversal curse phenomenon, i.e., training on ”A is B” does not make the model automatically generalize to ”B is A”. This failure again illustrates that generalization mechanisms taken for granted in humans do not necessarily occur in LLMs.
@@ 7 Discussion
We believe the main reason for this generalization gap is that past tense examples are out-of-distribution compared to the refusal examples used for fine-tuning, and current alignment techniques do not automatically generalize to them. Indeed, as we have shown in Section 4, correctly refusing on past tense examples is feasible via direct fine-tuning, and some models—like Llama-3 with the refusal-enhancing system prompt—are already relatively robust. Moreover, there are also other possible solutions that do not rely on SFT or RLHF, such as output-based detectors (Jain et al., 2023; Inan et al., 2023) and representation-based methods, including harmfulness probing (Ousidhoum et al., 2021) and representation rerouting (Zou et al., 2024). These approaches can reject harmful outputs, which seems to be an easier task compared to patching all possible inputs that can lead to harmful generations.
@@ Acknowledgements
We thank Francesco Croce for many insightful discussions and feedback on a draft of this paper. We are grateful to OpenAI for providing us API credits as part of the OpenAI Researcher Access Program. We thank Sudarsh Kunnavakkam for providing the API key to run the experiments on Claude-3.5 Sonnet. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship. This work was partially funded by an unrestricted gift from Google and by the Swiss National Science Foundation (grant number 212111).
@@ Appendix A Additional Details
Table 9: The system prompt for R2D2 from HarmBench [Mazeika et al., 2024].
@@
Table 11: The system prompt for GPT-4 as a semantic judge introduced in Chao et al. [2023].
@@
Table 12: The prompt for Llama-3-70B as a jailbreak judge from JailbreakBench [Chao et al., 2024].
@@
Table 13: The prompt for Llama-3 8B as a refusal judge from JailbreakBench [Chao et al., 2024].